Calculate the molarity of a KF (aq) solution containing 116.2 g of KF in 3.00 L of solution? [1] 0.02585 M

[2] 0.667 M

[3] 38.7 M

[4] 58.1 M

[5] None of the above


1
Expert's answer
2022-05-06T13:19:03-0400

M(KF)=39+19=58 g/mol

n(KF)=116.2/58=2 mol

c(KF)=n(KF)/Vsolution=2/3=0.667 mol/l

Answer: 1. 0.667 mol/l